# Are We Solving a Well-Defined Problem? A Task-Centric Perspective on Recommendation Tasks

### v2

@@ 3.3. Yet Another Rewritten Task Formulation
Now, we define the recommendation task as making a recommendation at time $t$ for user $u$ by referencing user-item interactions that occurred before $t$ and selecting suitable candidate items that satisfy the selection criteria $s(\cdot)$ to meet the inferred user preference. Following the example of repeat consumption and exploration, the final recommendation items presented to a user could be the results of running multiple recommendations made from multiple sets of candidate items conditioned on different multiple $s(\cdot)^{\prime}s$. Nevertheless, in terms of model evaluation, it is more meaningful to independently evaluate each model specific to candidate items on one selection function $s(\cdot)$.
@@ 4. From Recommendation to User Consumption
Figure 4. User costs after recommendations made at time $t$: from pre-interaction judgment, to interaction, then to an optional post-interaction feedback. Among the recommended items, $i_{2}$ satisfies selection criteria $S(\cdot)$.
So far, we have discussed the importance of explicitly incorporating the time factor into the problem definition and examined suitable candidate items for recommendation based on the context of the recommendation task. We consider these to be essential factors in the task definition due to their direct impact on model evaluation, which in turn directly impact the model design. Next, we take a closer look at user-item interactions $U\times I$ and analyze their meaning, as instances $(u,i,t_{x})$ are widely used to train recommender systems in academic research.
There are many reasons for a user to interact with an item, such as through an intentional search rather than a recommendation. Since we are discussing recommendations, we make a general assumption that if a user interacts with an item, it reflects some form of preference for that item. However, we aim to explore the various stages from a recommended item $i\in R_{t}^{u}$ to an interaction $(u,i,t_{x})$, and to what extent a model can learn from the interaction. This process can be considered as another form a mapping:
@@
4.1. User-Item Interaction: The Life Cycle
In the early stages of RecSys research, many studies focused on explicit user feedback, such as ratings e.g., a user giving a movie a 5 out of 5 rating (Herlocker et al., 2004). However, since explicit feedback is often difficult to obtain, implicit feedback has become the mainstream in academic research e.g., a user has watched a movie. Nevertheless, Ricci et al. (2011) suggest that user preference can also be inferred from actions such as navigating to a product page e.g., a user clicking on a movie for more details among a list of candidate movies. However, there is a lack of detailed discussion on to what extent such feedback contribute to accurately modeling user preference.
In our discussion, we divide user-item interactions into three stages: pre-interaction judgment, interaction, and post-interaction feedback, illustrated in Figure 4. While these stages may not apply to all recommendation scenarios, they provide a useful framework for understanding the relationship between user-item interactions and user preferences.
Consider a user searching for a hotel in a city she has never visited, for a conference. She is presented with a list of 10 hotels on a booking website, initiating the pre-interaction judgment stage. She clicks on one hotel for more details, a decision influenced by factors such as images, branding, location, price, or other readily available attributes. Next, she enters the deliberate evaluation phase, where she reads the hotel description, checks room facilities, and reviews user feedback before deciding to book. All these effort happens before the user stays at the hotel.
A few months later, she attends the conference and stays at the hotel. This marks the interaction stage, where she gains firsthand experience. Finally, after checking out, she provides feedback in the form of a rating and review, entering the post-interaction feedback stage. This feedback may consider various factors in a typical hotel review, such as location convenience, price, food, service quality, and hygiene.
Both quick judgments and deliberate evaluations in the pre-interaction stage come at a cost, the user’s time and effort in gathering relevant information. The outcomes of these efforts strongly indicate user preference, especially in similar future situations, e.g., planning another trip or booking a restaurant.
@@ 4.1. User-Item Interaction: The Life Cycle
Post-interaction feedback, however, may not always accurately reflect user preference. While a review can highlight the primary reasons for choosing a hotel, representing genuine user preference, it may also describe aspects such as expectation gaps and booking effort, which are not necessarily preference indicators. A rating, influenced by these factors, may not always be a reliable measure of preference compared to the concrete fact that the user chose to stay at the hotel. In fact, in many practical scenarios, post-interaction feedback is often difficult to collect, as it requires users to put in additional effort to provide a rating or write a review. Hence, in our following discussion, we focus on the other two stages.
In general, recommendations are presented as a list of candidate items before any user-item interaction occurs. At this stage, the user relies solely on their judgment, which inherently reflects preference, as they would likely make the same choice in similar circumstances. However, most academic datasets only contain actual interactions, lacking context on the reasons behind them. This limitation makes it challenging to model user preference purely based on observed interactions. Thus, a deeper understanding of the user decision-making process is crucial for more effective preference modeling.
@@ 4.2. Complexity of Pre-Interaction Judgment
We now focus on the pre-interaction judgment, which often takes the form of quick judgment or deliberate evaluation before a choice is made. The complexity of this judgment from the user’s perspective may arise from several dimensions.
Informed vs Uninformed Decision. One major factor from the user’s perspective is whether they possess the knowledge to accurately judge an item before interacting with it. For familiar items like books, movies, or other products the user has prior experience with, they can make an informed decision based on available attributes or other relevant information. Take movies as an example, users may decide whether to watch a film based on the director, cast, genre, a brief synopsis, or simply the poster. However, if a user has never used a robot vacuum before, many of the terms in the product description may be unfamiliar to them. Even after reading the machine specifications and user reviews, they may struggle to discern the pros and cons of a specific model, with references to her own home layout and floor conditions.
These uninformed decisions may not necessarily indicate user preferences. It is also challenging for a recommendation platform to determine whether a user’s decision was based on prior knowledge or made without a full understanding of the item. This distinction is crucial, as recommendations based on uninformed interactions might not truly reflect user interests, potentially leading to less effective personalization. For example, after using a robot vacuum for some time, the user may realize that he should purchase another model with features better suited to his floor conditions.
Items in One vs. Multiple Types/Categories. Many studies in recommendation are conducted on datasets containing only one type of item, such as books, movies, news, or music. In these cases, there are common characteristics, such as genre, director, or artist, that users can rely on for pre-interaction judgment before actually interacting with a recommended item. However, in the context of online shopping, where products span thousands of categories, users apply different criteria and expectations when making judgments for different types of items, for informed decisions.
@@ 4.3. Recognition of User-Item Interaction
After the pre-interaction judgment phase, the user makes a decision and interacts with an item. In most datasets used for RecSys evaluation, user-item interactions are recorded in the form of $(u,i,t_{x})$. However, interactions may appear in different forms depending on the recommendation context.
@@
Interaction Process Can Be Complicated. Taking online shopping as an example, the process does not end when a user clicks on a recommended item. After deciding on a product, the user might add the item to their cart, proceed to payment, receive the delivery, and ultimately complete the purchase. From the shopping platform’s perspective, this sequence of events marks a successful interaction. However, complications can arise if the user later decides to return the product due to reasons such as quality issues, unmet expectations, or simply a change of mind.
@@
This raises an important question: should an unsuccessful purchase (i.e., one that results in a return) still be considered a valid user-item interaction for learning user preferences? On one hand, the initial decision to purchase the item indicates some level of interest or preference. On the other hand, the return suggests dissatisfaction or misalignment with the user’s expectations. If returns are not accounted for, the model might incorrectly reinforce recommendations for similar items, leading to suboptimal suggestions. Therefore, when incorporating interaction data into preference modeling, it is crucial to differentiate between successful and unsuccessful purchases and consider additional contextual signals, such as return rates, to better understand user preferences.
@@
Absence of Pre-Interaction Judgment. Some of the user-item interactions happen without a pre-interaction judgment phase. This is especially common in scenarios like music streaming and short-video viewing, where users often do not actively select each item. Instead, they are presented with an initial set of options, and after selecting the first item, subsequent content is automatically fed to users by the recommendation engine, e.g., a playlist or streaming.
@@
In such cases, user engagement signals, such as skipping, fast-forwarding, or continuing to watch/listen, play a crucial role in modeling preferences. Unlike traditional recommendation settings where users consciously evaluate and select items before interaction, here, user preferences are inferred more dynamically based on real-time behaviors, or the interaction itself. In fact, users are less likely to explicitly express “like” in a streaming setting but are more likely to express “dislike”. This distinction raises important challenges for recommendation models, as they must distinguish between passive exposure and active preference while adapting to continuously evolving user interests. In this case, the common form of user-item interaction $(u,i,t_{x})$ becomes less accurate compared to other settings. Such recommendation in streaming form also post questions in item attribute modeling e.g., evaluating whether the cover image of a short video influence user viewing as the user may not even has the chance of viewing the cover image for each video in the streaming.
@@
Unobservable Interaction. There are also recommendation scenarios where user-item interactions are not directly observable and can only be inferred from external sources.
@@ 4.4. Interdependency Across Recommendations
Recommendations can occur either independently, as in hotel booking, or within a session-based context, such as music streaming and short-video viewing. In the latter case, subsequent recommendations may be influenced by the previous selections or even the initial choice made at the start of the session.
Figure 5. Reasons for a bad recommendation. The option “unrelated to the current video” suggests the dependency between the next recommendation and the current item.
Figure 5 shows a feedback form for assessing video recommendation quality from YouTube. Notably, one of the options for a poor recommendation is labeled as “unrelated to the current video.” This indicates that the recommendation system tailors the next suggested video based on the one currently being played. The inclusion of this feedback option emphasizes users’ expectations that recommendations within a session should maintain contextual relevance. This is a key characteristic of session-based recommendation, a specialized type of recommendation task (Li et al., 2025). A detailed discussion on the recommendation flow is made in (Sun, 2024).
This example further highlights the complexity of recommendation systems and the variety of factors that need to be accounted for when designing different recommendation tasks. Each recommendation scenario, whether based on user preferences, session context, or product types, requires careful formulation to ensure that the user experience is optimized and that the system accurately reflects user intent.
@@ 5. Discussion and Perspectives
Next, we discuss the ultimate goal of recommendation, i.e., reducing user cost, the relationship between task formulation, solution, and evaluation, and the intersection of task specificity and model generalizability, along with our perspectives on RecSys.
@@ 5.1. Recommendation vs User Cost
The ultimate goal of a recommender system is to reduce user effort in finding products or services of interest, enhance their enjoyment of recommendations, and build trust in the system. However, users still incur different costs at various stages of the interaction process. In the pre-interaction judgment stage, users must evaluate recommendations based on available information, such as descriptions, images, reviews, or other metadata, which requires cognitive effort and time. This cost is higher when users lack prior knowledge of the item category or when the recommendation list is overwhelming. During the interaction stage, users experience the actual product or service, which may involve monetary costs (e.g., purchasing a product), time investment (e.g., watching a recommended movie), or engagement effort (e.g., exploring an unfamiliar interface). If the recommendation is poor, users may feel frustrated, leading to dissatisfaction and disengagement. Notably, such negative experiences can accumulate over time, eventually causing users to leave the platform, confirmed by a recent large scale survey (Zou et al., 2024). Finally, in the post-interaction stage, users may be asked to provide feedback, such as ratings or reviews. While valuable for possibly improving recommendation quality, this step requires additional effort, and many users may choose not to participate.
Understanding and minimizing all such costs at each of the three stages is essential for enhancing user experience and optimizing the effectiveness (and even trust) of recommender systems. The task definition discussed above primarily expresses the inputs that a recommender model should consider and the desired output. However, the various costs incurred by users at different stages are not explicitly represented in the problem definition, but can be incorporated into the loss function design in the proposed models.
Table 1. Example recommendation tasks across different domains. The characteristics are based on our best understanding, which may not be applicable to any specific platform. The “price” attribute is ignored from pre-interaction judgment.
Domain Item Type Session Candidate Item Selection Pre-Interaction Judgment User-Item Interaction
Music streaming Single Yes Repeat, Exploration No, except the first song Progress ratio, Skip
Short video Single Yes No, or non-repeat No, except the first video Progress ratio, Watching speed, Skip
Online shopping Multi Might Repeat, Exploration Description, Image, Review Payment, Return
News Single Yes Region, Genre Title, Source Click, View
Job Single No Location Description Unobservable, Post-interaction inference
Course Single No Meeting pre-requisite Description, Instructor, Institute Progress, completion
Food delivery Single No Repeat, Exploration, Delivery distance/time Image, Description, Review Payment
POI Multi Might Region, Repeat, Exploration Description, Image, Travel cost, Review Unobservable, Post-interaction inference
@@ 5.2. Task, Solution, and Evaluation
Some existing RecSys tasks are not well formulated. Sequential recommendation and intent-aware recommendation are two examples; they do not fundamentally differ from a typical recommendation problem. In the survey paper, Jannach and Zanker (2024) defines: intent-aware recommender “is a recommender system that is designed to capture the users’ underlying current motivations and goals in order to support them.” If we view the problem definition by its inputs and outputs, there is no significant difference from the definition in Equation 6. Probably due to a less distinctive problem formulation, there are questions on the progress made (Shehzad et al., 2025). In the survey paper (Pan et al., 2024), the task of sequential recommendation is defined as follows: “In sequential recommendation, we often have one or more sequences of interacted items w.r.t. each user, as well as some auxiliary information to help learn user preferences. Our goal is then to generate a ranked list of items accurately for each user”. This is basically a different view of the very same $U\times I$. Naturally all items interacted with by a user form a sequence as illustrated in Figure 2. Solutions that pay attention on such sequences may be able to achieve a better recommendation accuracy, particularly in the RecSys settings with items of a single type e.g., music, movie, book, and news. However, the task to be addressed remains a generic recommendation.
A comparable research challenge in the past was imbalanced classification. In that domain, all classification problems with imbalanced class distributions were broadly categorized as “imbalanced classification problems,” and researchers sought a common solution to address the imbalance while often overlooking the true underlying reasons for it. However, data imbalance can arise from multiple causes, and while imbalance is an easily observable characteristic of data distribution, it is not itself the root cause (López et al., 2013). As a result, solutions designed specifically for “imbalanced classification”, the surface form of the data imbalance, may or may not be effective, leading to inconclusive results.
The proposed task-centric framework provides a set of dimensions to consider and compare RecSys tasks. Accordingly, the proposed solutions can be evaluated in a fair and meaningful way.
@@ 5.3. Task Specificity vs Model Generalizability
One potential approach to addressing this challenge is to establish a broad categorization of recommendation tasks based on key dimensions, such as candidate item selection, single- vs. multi-category recommendations, the cost and information available during pre-interaction judgment, the nature of user-item interactions, and the availability of post-interaction feedback. Table 1 provides an overview of several example recommendation domains and highlights their distinguishing characteristics.
From the table, we observe that music streaming and short-video recommendations share many common characteristics, except that users rarely prefer to repeatedly watch the same short videos. However, both domains differ significantly from POI (Point of Interest) recommendation. In POI recommendation, items include restaurants, parks, and historical landmarks, making it a multi-category recommendation task. Additionally, there is a fundamental distinction between recommending previously visited places (e.g., on a map or navigation interface) and suggesting new locations for exploration. This contrast underscores the importance of defining recommendation tasks based on their inherent characteristics rather than treating them as a single unified problem.
Such categorization ensures that all solutions are fairly evaluated on benchmark datasets regardless of the techniques used. We believe that, a detailed understanding and categorization of research tasks can provide insights into selecting appropriate baselines (Ivanova et al., 2023) and datasets (Beel et al., 2024). Ultimately, the effectiveness of a recommender system depends not necessarily on the sophistication of the model but on how well it aligns with the specific characteristics of the recommendation task at hand. Striking the right balance between generalizability and task specificity remains an open challenge and a key research direction for advancing the field.
By categorizing recommendation tasks along these dimensions, we provide a structured framework for understanding the diversity of recommender system applications. Future work should focus on refining these categorizations and developing standardized evaluation protocols that account for the unique characteristics of each recommendation task.

### v3

@@ 3.3. Task Formulation with Constraints
The final items recommended to a user could be the results of running multiple recommendations made from multiple sets of candidate items conditioned on different multiple $s(\cdot)$’s. In terms of model evaluation, it is more meaningful to independently evaluate each model specific to candidate items on one selection function $s(\cdot)$.
@@ 4. From Recommendation to User Consumption
Users may interact with items for various reasons. An interaction could be the result of an intentional search rather than recommendation. For instance, the Yambda-5B dataset includes an is_organic flag to indicate actions not driven by recommendations (Ploshkin et al., 2025). In this discussion, we generally assume that a user’s interaction with an item reflects some degree of preference for it.
Our focus is to examine the stages from a recommended item $i\in R_{t}^{u}$ to an interaction $(u,i,t_{x})$, and to what extent a model can learn from such interactions. This process can be considered as another form a mapping:
@@
4.1. The Life Cycle of User-Item Interaction
We divide user-item interactions into three stages: pre-interaction judgment, interaction, and post-interaction feedback. These stages may not apply to all recommendation scenarios, but offer a useful framework for understanding the relationship between user interactions and preferences. A conceptually similar framework was proposed by Lee and Kim (2023), who distinguish between pre-use preference (a user’s impression of an item before interacting with it) and post-use preference (the impression formed after interaction). Our discussion aims to highlight the different types of pre-interaction and interaction across various recommendation scenarios, which differs from both the technical solutions presented in (Lee and Kim, 2023) and the decision-making perspectives discussed in (Jameson et al., 2022).
Consider a user booking a hotel in an unfamiliar city for a conference. When she first browses a list of options, she forms a pre-interaction judgment based on visible cues such as images, branding, location, and price. After clicking on one hotel, she enters the deliberate evaluation stage — reading descriptions, checking facilities, and reviewing feedback before booking. Months later, her stay in the hotel marks the interaction stage. After checkout, she leaves a rating and review, marking the post-interaction feedback stage.
@@ 4.1. The Life Cycle of User-Item Interaction
Both quick judgments and deliberate evaluations in the pre-interaction stage come at a cost, the user’s time and effort in gathering relevant information. The outcomes of these efforts typically indicate user preference, especially in similar future situations, e.g., planning another trip. Post-interaction feedback, however, may not always accurately reflect user preference. While a review can highlight the primary reasons for choosing a hotel, representing genuine user preference, it may also describe aspects such as expectation gaps and booking effort, which are not necessarily preference indicators. A rating, influenced by these factors, may not always be a reliable measure of preference compared to the fact that the user chose to stay at the hotel. In many practical scenarios, post-interaction feedback is often difficult to collect, as it requires users to put in additional effort. Hence, in our following discussion, we focus on the other two stages.
@@ 4.2. Complexity of Pre-Interaction Judgment
The complexity of pre-interaction judgment from the user’s perspective may arise from several dimensions.
4.2.1. Informed vs Uninformed Decision
One major factor from the user’s perspective is whether they possess the knowledge to accurately judge an item before interacting with it. For familiar items like books, movies, or other products the user has prior experience with, they can make an informed decision based on available attributes or other relevant information. Take movies as an example, users may decide whether to watch a film based on the director, cast, genre, a brief synopsis, or simply the poster. However, if a user has never used a robot vacuum before, many of the terms in the product description may be unfamiliar to them. Even after reading the machine specifications and user reviews, she may struggle to discern the pros and cons of a specific model, with references to her own home layout and floor conditions.
Uninformed decisions may not necessarily indicate user preferences. It is also challenging for a recommendation platform to determine whether a user’s decision was based on prior knowledge or made without a full understanding of the item. This distinction is crucial, as recommendations based on uninformed interactions might not truly reflect user interests, potentially leading to less effective personalization. For example, after using a robot vacuum for some time, the user may realize that she should purchase another model with features better suited to her floor conditions.
4.2.2. Items in One vs. Multiple Types
Many studies in recommendation are conducted on datasets containing only one type of item, such as books, movies, news, or music. In these cases, there are common characteristics, such as genre, director, or artist, that users can rely on for pre-interaction judgment before actually interacting with a recommended item. However, in the context of online shopping, where products span thousands of categories, users apply different criteria and expectations when making judgments for different types of items.
@@ 4.3. Recognition of User-Item Interaction
User-item interactions are often recorded in the form of $(u,i,t_{x})$ in a RecSys dataset. However, interactions may appear in different forms depending on the recommendation context.
4.3.1. Interaction Process Can Be Complicated.
@@
Taking online shopping as an example, the process does not end when a user clicks on a recommended item. After deciding on a product, the user might add the item to their cart, proceed to payment, receive the delivery, and ultimately complete the purchase. From the shopping platform’s perspective, this sequence of events marks a successful interaction. However, complications can arise if the user later decides to return the product due to reasons such as quality issues, unmet expectations, or simply a change of mind.
@@
This raises an important question: should an unsuccessful purchase (i.e., one that results in a return) still be considered a valid user-item interaction for learning user preferences? On the one hand, the initial decision to purchase the item indicates some level of interest or preference. On the other hand, the return suggests dissatisfaction or misalignment with the user’s expectations. If returns are not accounted for, the model might incorrectly reinforce recommendations for similar items, leading to suboptimal suggestions. Therefore, when incorporating interaction data into preference modeling, it is crucial to differentiate between successful and unsuccessful purchases and consider additional contextual signals, such as return rates, to better understand user preferences.
4.3.2. Interaction without Pre-Interaction Judgment.
@@
Some user-item interactions happen without a pre-interaction judgment phase. This is especially common in scenarios like music streaming and short-video viewing, where users often do not actively select each item. Instead, they are presented with an initial set of options, and after selecting the first item, subsequent content is automatically fed to users by the recommendation engine, e.g., a playlist or streaming.
@@
In such cases, user engagement signals, such as skipping, fast-forwarding, or continuing to watch/listen, play a crucial role in modeling preferences. Unlike traditional recommendation settings where users consciously evaluate and select items before interaction, here, user preferences are inferred more dynamically based on real-time behaviors, or the interaction itself. Recommendation models must distinguish between passive exposure and active preference while adapting to continuously evolving user interests. In this case, the common form of user-item interaction $(u,i,t_{x})$ becomes less accurate compared to other settings. Such recommendation in streaming form also post questions in item attribute modeling e.g., evaluating whether the cover image of a short video influence user viewing as the user may not even has the chance of viewing the cover image for each video in the streaming.
4.3.3. Unobservable Interaction.
@@
There are also recommendation scenarios where user-item interactions are not directly observable and can only be inferred from external sources.
@@ 4.4. Interdependency across Recommendations
Recommendations can occur either independently, as in hotel booking, or within a session-based context, such as music streaming and short-video viewing. In the latter case, subsequent recommendations may be influenced by the previous selections or even the initial choice made at the start of the session. This is a key characteristic of session-based recommendation, a specialized type of recommendation task (Li et al., 2025). A detailed discussion on the recommendation flow is made in (Sun, 2024).
Each recommendation scenario, whether based on user preferences, session context, or product types, requires careful formulation to ensure that the user experience is optimized and that the system accurately reflects user intent.
@@ 5. Discussion and Perspectives
Next, we discuss the ultimate goal of recommendation, examine the relationships among task formulation, solution, and evaluation, explore the intersection of task specificity and model generalizability, and provide actionable guidance for RecSys academic research.
@@ 5.1. Recommendation vs User Cost
The ultimate goal of a recommender system is to reduce user effort in finding products or services of interest, enhance their enjoyment of recommendations, and build trust in the system. However, users still incur different costs at various stages of the interaction process. In the pre-interaction judgment stage, users evaluate recommendations based on available information, such as descriptions, images, reviews, or other metadata, which requires cognitive effort and time. During the interaction stage, users experience the actual product or service, which may involve monetary costs (e.g., purchasing a product), time investment (e.g., watching a recommended movie), or engagement effort (e.g., exploring an unfamiliar interface). If the recommendation is poor, users may feel frustrated, leading to dissatisfaction and disengagement (Zou et al., 2024). Finally, in the post-interaction stage, users may be asked to provide feedback, such as ratings or reviews. This step requires additional effort, and many users may choose not to participate.
Understanding and minimizing these costs at each of the three stages is essential for improving user experience and optimizing the effectiveness, and even the trustworthiness, of recommender systems. However, as illustrated in the examples above, such costs manifest differently across recommendation scenarios. The RecSys task definition in Equation 6 primarily specifies the inputs a recommender model should consider and the desired output. Yet, the various costs incurred by users at different stages from recommendation to interaction (Equation 7) cannot be explicitly represented in a formal formulation. Nevertheless, understanding these costs can inform the design of more refined loss functions for learning recommendation models. For example, in short-video recommendation, videos that a user watches for a duration shorter than their average viewing percentage can be treated as tolerance samples. In (Zou et al., 2024), such samples are assigned different losses compared to fully watched videos, leading to improvements in user retention verified through A/B testing. In this context, watching an uninteresting video constitutes a cost to the user.
@@ 5.2. Task, Solution, and Evaluation
Some existing RecSys tasks are not well formulated. Sequential recommendation and intent-aware recommendation are two examples; they do not fundamentally differ from a typical recommendation problem. In the survey paper, Jannach and Zanker (2024) defines: intent-aware recommender “is a recommender system that is designed to capture the users’ underlying current motivations and goals in order to support them.” If we view the problem definition by its inputs and outputs, there is no significant difference from the definition in Equation 6. Probably due to a less distinctive problem formulation, there are questions on the progress made (Shehzad et al., 2025). In the survey paper (Pan et al., 2026), the task of sequential recommendation is defined as follows: “In sequential recommendation, we often have one or more sequences of interacted items w.r.t. each user, as well as some auxiliary information to help learn user preferences. Our goal is then to generate a ranked list of items accurately for each user”. This is basically a different view of the very same $U\times I$. Naturally all items interacted with by a user form a sequence as illustrated in Figure 1. Solutions that pay attention on such sequences may be able to achieve a better recommendation accuracy, particularly in the RecSys settings with items of a single type e.g., music, movie, book, and news. However, the task to be addressed remains a generic recommendation.
@@ 5.3. Task Specificity vs Model Generalizability
5.4. From Theory to Practice
Moving forward, a practical step toward addressing the challenges discussed earlier is to establish a clear categorization of recommendation tasks. In our recent survey (Zou and Sun, 2025), which includes only research papers reporting online A/B testing results from production environments, we classify real-world recommendation tasks into two main categories. Transaction-oriented recommender systems generate item recommendations with the primary goal of prompting transactional user actions, optimizing for metrics such as conversion rate, revenue, or purchase likelihood. E-commerce platforms are typical examples. Content-oriented recommender systems, on the other hand, generate recommendations to facilitate user consumption and engagement, optimizing for metrics such as dwell time, clicks, or user satisfaction. Common examples include news, video, and music recommenders. Although these two main categories do not account for all types of recommendation systems, they capture the majority of widely studied recommendation scenarios.
The objectives of recommender systems across these two broad categories differ substantially and can vary further within each category. Consequently, more specific recommendation tasks can be refined along key dimensions discussed earlier. For instance, recency is a key factor in news recommendation but may be less relevant for other types of content-oriented recommendation. RecSys research focused on algorithmic advancement should clearly specify the recommendation task it addresses, use datasets and evaluation settings that reflect realistic conditions, and compare against baselines designed for similar contexts.
Such categorization ensures fair and meaningful evaluation across different recommendation tasks. Importantly, this emphasis on task specificity also requires a shared understanding within the community, especially among reviewers, that researchers should not be expected to overgeneralize their solutions to unrelated datasets or problem settings. For example, a submission should not be penalized for not reporting results on widely used datasets like MovieLens, when it addresses fundamentally different recommendation scenarios. The effectiveness of a recommender system depends much on how well it aligns with the specific characteristics of the target task.
We highlight two recent examples that suggest researchers and reviewers are increasingly moving in this direction. Charolois-Pasqua et al. (2025) propose a playlist-generation recommender focused on a specific task: generating playlists from titles. Their method is evaluated on the Million Playlist Dataset and compared against task-relevant baselines. The authors further analyze user effort in playlist creation and assess the usefulness of playlist titles, demonstrating strong task alignment rather than pursuing cross-domain generalization. Similarly, based on their experience building a small-scale production news recommender system, Higley et al. (2025) observe that “published models are surprisingly difficult to apply to the kinds of data found in real-world datasets and practical recommendation problems.” They emphasize that “building recommender systems that serve real users requires deep and specific engagement not just with a domain in general, but with the particular characteristics of specific datasets, applications, and user communities,” underscoring the task-specific nature of RecSys research. Task-specific RecSys research has become increasingly feasible thanks to dataset availability. For example, Yambda-5B includes both implicit (e.g., listening) and explicit (e.g., likes and dislikes) feedback, as well as organic user actions specific to the music streaming recommendation task (Ploshkin et al., 2025). The authors also introduce a global temporal split to better reflect real-world settings and prevent data leakage.